chemistry

The reaction of water and coal at a high temperature produces a mixture of hydrogen and carbon monoxide gases. This mixture is known as synthesis gas (or syngas). Whatmass of hydrogen gas can be formed from the reaction of 51.3 g of carbon with excess water?

C(s) + H2O(g) ? H2(g) + CO(g)
Answer #1

The balanced equation is

C + H2O = CO + H2

moles C = 71.3 g /12.011 g/mol= 5.94
the ratio between C and CO is 1 : 1

moles CO = 5.94

mass CO = 5.94 mol x 28.01 g/mol=166.4 g

answered by: tod